Interpret the meaning of "The 95% confidence interval for true mean forced vital capacity (FVC) is 4.56 +/or- 0.35."

Choose one.

a. The true population mean forced vital capacity (FVC) is between 4.21 and 4.91

b. We are 95% confident that the true population mean forced vital capacity (FVC) is between 4.21 and 4.91. in the long run, a confidence interval constructed this way will capture the true value of mean 95% of the time

c. In the long run, a confidence interval constructed this way will capture the true value of mean 99% of the time

d. The margin of error = 0.96
Answer #1

Lower limit = 4.56-0.35 = 4.21

Upper limit = 4.56+0.35 = 4.91

Ans:

b. We are 95% confident that the true population mean forced vital capacity (FVC) is between 4.21 and 4.91. in the long run, a confidence interval constructed this way will capture the true value of mean 95% of the time